OpenStudy (anonymous):

Show that the arclength of the curve y=(4-x^2/3)^3/2 for x=0 to x=8, as illustrated, is an improper integral and evaluate it.

OpenStudy (dumbcow):

arc length \[\int\limits_0^8 \sqrt{1+(\frac{dy}{dx})^{2}} dx\]

OpenStudy (dumbcow):

\[\frac{dy}{dx} = -\frac{\sqrt{4-x^{2/3}}}{x^{1/3}}\]

OpenStudy (dumbcow):

i dont see how this becomes an improper integral?

myininaya (myininaya):

The function we are integrating wouldn't be continuous at x=0 so the integral is improper.